OpenStudy (jenniferjuice):

Which of the following is not a way to represent the solution of the inequality 8x − (5x + 4) ≥ −31?

OpenStudy (jenniferjuice):

x ≥ −9 −9 ≤ x A number line with a closed circle on −9 and shading to the left. A number line with a closed circle on −9 and shading to the right.

OpenStudy (anonymous):

8x - (5x + 4) >= -31 8x -5x - 4 >= -31 3x - 4 >= -31 3x >= -31 + 4 3x >= - 27 x >= -9 closed circle on -9 shaded to the left is INCORRECT